Let V be the real vector space of polynomials
over R of degree at most 2 and W be the
subspace of V generated by (1+x², 1+2x² }.
Find the kernel of the differential linear
operator d/dx on W
1
Expert's answer
2020-12-20T14:51:16-0500

"V=\{a_{0}+a_{1}x+a_{2}x^{2}|a_{0},a_{1},a_{2}\in R\}\\
\\
W=<1+x^{2},1+2x^{2}>"

Any vector of W is of the form "a(1+x^{2})+b(1+2x^{2})" =(a+b)+(a+2b)"x^{2}"

Let T="\dfrac{d}{dx}"

As T operates on W, according to the definition of differential operator

T("(a+b)+(a+2b)x^{2})=2x(a+2b)"

Hence Kernel (T)="\{p\in W|p=0\}" where p is a polynomial of W.

"\Rightarrow T ((a+b)+(a+2b)x^{2}=0"

"\Rightarrow 2x(a+2b)=0"

"\Rightarrow (a+2b)=0"

"\Rightarrow a=-2b"


Therefore P=(-2b+b)+(-2b+2b)x"^{2}" =-b , which is a constant polynomial.


As "b\in R\\" , Kernel(T) is isomorphic to R.